OpenStudy (anonymous):

What is the total sample size (in grams) for a sample of magensium nitrite which contains 1.09 g of oxygen?

OpenStudy (mos1635):

Mg(NO3)2 1 molucule of Mg(NO3)2 contains 6 atoms Of Oxygen 1 mole of Mg(NO3)2 contains 6 moles Of Oxygen 1mole of Mg(NO3)2 is 24+2*(14+3*16)=148 gr 6 mole of atoms of oxygen is 6*16=96gr so 148 gr of Mg(NO3)2 contains 96gr Of Oxygen x=? 1.09

OpenStudy (vincent-lyon.fr):

Hi! Do you actually mean magnesium nitrite, or is your problem about magnesium nitrAte?

OpenStudy (anonymous):

i mean magnesium nitrite

OpenStudy (anonymous):

I'm always a fan of starting off with a balanced chemical equation

OpenStudy (vincent-lyon.fr):

@zbay: there is no reaction involved in this problem. @ktrigga: mos1635's method is correct, but the nitrite ion's formula is\[(NO _{2})^{-}\] , so there are only 2 atoms of oxygen and not 3.

OpenStudy (mos1635):

Mg(NO4)2 1 molucule of Mg(NO2)2 contains 4 atoms Of Oxygen 1 mole of Mg(NO2)2 contains 4 moles Of Oxygen 1mole of Mg(NO2)2 is 24+2*(14+2*16)=124 gr 4 mole of atoms of oxygen is 4*16=64gr so 124gr of Mg(NO3)2 contains 96gr Of Oxygen x=? 1.09 thanks @Vincent-Lyon.Fr